# Equal contribution

### v1

@@ Abstract
We propose a variational inference approach to sample from the posterior distribution for solving inverse problems. From a pre-trained diffusion model, our approach trains a conditional flow model to minimize the divergence between the proposal variational distribution and the posterior distribution implicitly defined through the diffusion model. Once trained, the flow model is capable of sampling from the posterior distribution with a single NFE, amortized with respect to the measurement. The proposed method paves a new path for distilling a diffusion prior for efficient posterior sampling. We show that our method is applicable to standard signals in Euclidean space, as well as signals on manifold.
@@ 1 Introduction
${\bm{y}}={\mathcal{A}}({\bm{x}})+{\bm{n}},\quad{\bm{y}}\in{\mathbb{R}}^{m},\,{\bm{x}}\in{\mathbb{R}}^{n},\,{\bm{A}}:{\mathbb{R}}^{n}\mapsto{\mathbb{R}}^{m},\,{\bm{n}}\sim{\mathcal{N}}(0,\sigma_{y}^{2}{\bm{I}}),$ (1)
where the goal is to infer an unknown signal ${\bm{x}}$ from the degraded measurement ${\bm{y}}$ obtained through some forward operator ${\mathcal{A}}$, leveraging the information contained in the measurement and the prior $p({\bm{x}})$. A powerful modern way to define the prior is through diffusion models [ho2020denoising, song2020score], where we train a parametrized model ${\bm{s}}_{\theta}$ to estimate the gradient of the log prior $\nabla_{\bm{x}}\log p({\bm{x}})$.
Solving inverse problems with the diffusion model can be achieved through posterior sampling with Bayesian inference. Arguably the standard way to achieve this is through modifying the reverse diffusion process of diffusion models. This adjustment shifts the focus from sampling from the trained prior distribution $p_{\theta}({\bm{x}}_{0})$ to sampling from the posterior distribution $p_{\theta}({\bm{x}}_{0}|{\bm{y}})$. This transition is facilitated by employing iterative projections to the measurement subspace [kadkhodaie2020solving, song2020score, chung2022come, wang2023zeroshot], guiding the samples through gradients pointing towards measurement consistency [chung2023diffusion, song2023pseudoinverseguided].
It should be noted that diffusion models learn the gradient of the prior and diffusion samplers [song2020denoising, lu2022dpm, song2020score] are methods that numerically solve the probability-flow ODE (PF-ODE) that defines the reverse diffusion sampling trajectory. Consequently, regardless of the specifics of the methods, standard diffusion model-based inverse problem solvers (DIS), even those that are considered fast, take at least a few tens of neural function evaluation (NFE), making them less effective for time-critical applications such as medical imaging and computational photography.
@@
Figure 2: Concept of the proposed method. (a) Training can be performed in an unsupervised fashion with a dataset consisting of degraded measurements ${\bm{y}}$ to train a conditional normalizing flow $G_{\phi}$ with the diffusion prior ${\bm{s}}_{\theta}$. (b) Once trained, one can achieve multiple posterior samples by inputting different noise vectors ${\bm{z}}\sim{\mathcal{N}}(0,{\bm{I}})$ concatenated with the condition ${\bm{y}}$ with a single NFE, generalizable across any measurement ${\bm{y}}$.
Another class of methods [feng2023score, feng2023efficient] introduces the use of variational inference (VI) to train a new proposal distribution $q_{\phi}^{\bm{y}}({\bm{x}})$ to distill the prior learned through the pre-trained diffusion model. The problem is defined as the following optimization problem
@@
where the superscript ${\bm{y}}$ emphasizes that the proposal distribution is specific for a single measurement ${\bm{y}}$. For tractable optimization, $q$ is often taken to be a normalizing flow [rezende2015variational, dinh2016density] (NF) so that the computation of likelihood can be done instantly, and one can sample multiple reconstructions from the posterior samples by plugging in different noise values from the reference distribution. Posing the problem this way yields a method that can achieve posterior samples with just a single NFE. Nevertheless, it is still impractical as training a measurement-specific variational distribution takes hours of training. It is often unrealistic to train a whole new model from scratch every time when a new measurement is taken.
In this work, we take a step towards a practical VI-based posterior sampler by distilling a diffusion model prior. To this end, we propose a conditional normalizing flow $q_{\phi}({\bm{x}}_{0}|{\bm{y}})$ as our variational distribution and amortize the optimization problem in Eq. (2) over the conditions ${\bm{y}}$. By using a network that additionally takes in the condition ${\bm{y}}$ as the input, we can train a single model that generalizes across the whole dataset without the need for cumbersome re-training for specific measurements. (See Fig. 2 for the conceptual illustration of the proposed method, as well as representative results presented in Fig. 1.) Interestingly, we find that the speed of optimization is not hampered with such amortization, and the proposed method achieves comparable performance against the measurement-specific flow model [feng2023efficient, feng2023score]. Furthermore, we extend the theory to consider inverse problems on the Riemannian manifold, showing that the proposed idea is generalizable even when the signal is not one the Euclidean manifold. In summary, our contributions and key takeaways are as follows
@@
    To the best of our knowledge, our method is the first diffusion prior distillation approach for solving inverse problems that are unsupervised (i.e. does not require any ground-truth data ${\bm{x}}$), as opposed to standard conditional NFs [lugmayr2020srflow] that required supervised paired data.
@@ 2.1 Score-based diffusion models
We adopt the standard framework for constructing a continuous diffusion process ${\bm{x}}(t)$, where $t\in[0,T]$ and ${\bm{x}}(t)\in\mathbb{R}^{d}$, as outlined by Song et al. [song2020score]. Specifically, our goal is to initialize ${\bm{x}}(0)$ from a distribution $p_{0}({\bm{x}})=p_{\text{data}}$, and evolve ${\bm{x}}(t)$ towards a reference distribution $p_{T}$ at time $T$, which is easily samplable.
@@
where $dt$ represents an infinitesimal negative time step, and $\bar{{\bm{w}}}$ denotes the backward standard Brownian motion. While the trajectory of Eq. (4) is stochastic, there also exists a corresponding probability-flow ODE (PF-ODE) that recovers the same law $p_{t}({\bm{x}})$ as the time progresses [song2020score, song2020denoising]
@@
This allows a deterministic mapping between the reference and the target distribution, and hence diffusion models can also be seen as a neural ODE [chen2018neural].
A neural network can be trained to approximate the true score function $\nabla_{\bm{x}}\log p_{t}({\bm{x}})$ through score matching techniques, as demonstrated in previous works [song2019generative, song2020score]. This approximation, denoted ${\bm{s}}_{\theta}({\bm{x}},t)\approx\nabla_{\bm{x}}\log p_{t}({\bm{x}})$, is then utilized to numerically integrate the reverse-time SDE. To effectively train the score function, denoising score matching (DSM) is often employed [hyvarinen2005estimation]
$\theta^{*}=\operatorname*{arg\,min}_{\theta}{\mathbb{E}}_{t\sim U(\varepsilon,1),{\bm{x}}(t),{\bm{x}}(0)}\left[\|{\bm{s}}_{\theta}({\bm{x}}(t),t)-\nabla_{{\bm{x}}_{t}}\log p({\bm{x}}(t)|{\bm{x}}(0))\|_{2}^{2}\right],$ (6)
Interestingly, the posterior mean, or the so-called denoised estimate can be computed via Tweedie’s formula [efron2011tweedie]. Specifically, for $p({\bm{x}}_{t}|{\bm{x}}_{0})={\mathcal{N}}({\bm{x}}_{t};\alpha_{t}{\bm{x}}_{0},\beta_{t}^{2}{\bm{I}})$,
@@ 2.2 Diffusion models for inverse problems (DIS)
While the former term can be replaced with a pre-trained diffusion model, the latter term is intractable and needs some form of approximation. Existing DIS [kawar2022denoising, chung2023diffusion, wang2023zeroshot] propose different approximations for $\nabla_{{\bm{x}}_{t}}\log p({\bm{y}}|{\bm{x}}_{t})$, which yields sampling from slightly different posteriors $\nabla_{{\bm{x}}_{t}}\log p_{\theta}({\bm{x}}_{t}|{\bm{y}})$.
Algorithmically, the posterior samplers are often implemented so that the original numerical solver for sampling from the prior distribution remains intact, while modifying the Tweedie denoised estimate at each time $\hat{\bm{x}}_{0|t}^{\theta}$ to satisfy the measurement condition given as Eq. (1). From Tweedie’s formula, we can see that this corresponds to approximating the conditional posterior mean ${\mathbb{E}}[{\bm{x}}_{0}|{\bm{x}}_{t},{\bm{y}}]$ in the place of the unconditional counterpart ${\mathbb{E}}[{\bm{x}}_{0}|{\bm{x}}_{t}]$. The algorithms are inherently iterative, and the modern solvers [chung2023diffusion, wang2023zeroshot, zhu2023denoising] require at least 50 NFE to yield a high-quality sample. Moreover, as existing methods can be interpreted as approximating the reverse distribution $p({\bm{x}}_{0}|{\bm{x}}_{t})$ with a simplistic Gaussian distribution $q({\bm{x}}_{0}|{\bm{x}}_{t})={\mathcal{N}}({\bm{x}}_{0};\hat{\bm{x}}_{0|t}^{\theta},s_{t}^{2}{\bm{I}})$ [peng2024improving], it typically yields a large approximation error, especially in the earlier steps of the reverse diffusion.
@@
Methods DIS Noise2Score [kim2021noise2score] RED-Diff Feng et al. [feng2023score, feng2023efficient] Ours
@@ 3.1 Variational inference in DIS
Standard DIS discussed in Sec. 2.2 sample from the posterior distribution by following the reverse diffusion trajectory. Another less studied approach uses VI to use a new proposal distribution, where the problem is cast as an optimization problem in Eq. (2). RED-diff [mardani2023variational] places a unimodal Gaussian distribution as the proposal distribution $q_{\phi}^{\bm{y}}({\bm{x}})$, and the KL minimization is done in a coarse-to-fine manner, similar to standard DIS, starting from high noise level to low noise level. While motivated differently, RED-diff and standard DIS have similar downsides of requiring at least a few tens of NFEs, as well as placing a simplistic proposal distribution. Furthermore, one can achieve only a single sample per optimization.
Recently, Feng et al. [feng2023score, feng2023efficient] uses an NF model for the proposal distribution while solving the same VI problem. The optimization problem involves computing the diffusion prior log likelihood $\log p_{\theta}({\bm{x}})$. It was shown that it can be exactly computed by solving the PF-ODE [feng2023score, song2020score], but numerically solving the PF-ODE per every optimization step is extremely computationally heavy, and hence does not scale well. To circumvent this issue, it was proposed to use a lower bound [feng2023efficient, song2021maximum]. Once trained, the NF model can be given different noise inputs ${\bm{z}}\sim{\mathcal{N}}(0,{\bm{I}})$ to generate diverse posterior samples with a single forward pass through the network. However, the training should be performed with respect to all the different measurements, not being able to generalize across the dataset. Our work follows along this path to overcome the current drawback as shown in Table 1.
@@ 3.2 Distillation of the diffusion prior
Our method involves distillation of the diffusion prior into a student deep neural network, in our case an NF model. Particularly, it involves evaluating the output of the model by checking the denoising loss gradients from the pre-trained diffusion model. This idea is closely related to variants of score distillation sampling (SDS) [poole2023dreamfusion, wang2024prolificdreamer], where the gradient from the denoising loss is used to distill the diffusion prior by discarding the score Jacobian. Possibly a closely related work is Diff-instruct [luo2024diff], where the authors propose to train a one-step generative model similar to GANs [goodfellow2014generative] by distillation of the diffusion prior with VI. By proposing an integral KL divergence (IKL) by considering KL minimization across multiple noise levels across the diffusion, it was shown that SDS-like gradients can be used to effectively train a new generative model. While having similarities, our method directly minimizes the KL divergence and does not require dropping the score Jacobian.
Orthogonal to the score distillation approaches, there have been recent efforts to train a student network to emulate the PF-ODE trajectory itself [song2023consistency, gu2023boot] with a single NFE, one of the most prominent directions being consistency distillation (CD) [song2023consistency]. While promising, the performance of CM is upper-bounded by the teacher PF-ODE. Thus, in order to leverage CD-type approaches for diffusion posterior sampling, one has to choose one of the approximations of DIS as its teacher model. In this regard, applying CD for diffusion inverse problem solving is inherently limited.
4 Proposed Method
@@ 4.1 Conditional NF for amortized score prior
The goal is to use a variational distribution that is conditioned on ${\bm{y}}$, such that the resulting distilled conditional NF model $G_{\phi}$ generalizes to any condition ${\bm{y}}$. To this end, inspired from the choices of [sun2021deep, feng2023score] we modify the objective in Eq. (2) to
@@
Computation of $\log p_{\theta}(\hat{\bm{x}}_{0})$ is more involved: to exactly compute the value, we would have to solve the PF-ODE, which is compute-heavy [song2020score, feng2023score]. To circumvent this burden, we leverage the evidence lower bound (ELBO) [song2021maximum, feng2023efficient] $b_{\theta}(\hat{\bm{x}}_{0})\leq\log p_{\theta}(\hat{\bm{x}}_{0})$:

### v2

# Amortized Posterior Sampling with Diffusion Prior Distillation

@@ Abstract
We propose Amortized Posterior Sampling (APS), a novel variational inference approach for efficient posterior sampling in inverse problems. Our method trains a conditional flow model to minimize the divergence between the variational distribution and the posterior distribution implicitly defined by the diffusion model. This results in a powerful, amortized sampler capable of generating diverse posterior samples with a single neural function evaluation, generalizing across various measurements. Unlike existing methods, our approach is unsupervised, requires no paired training data, and is applicable to both Euclidean and non-Euclidean domains. We demonstrate its effectiveness on a range of tasks, including image restoration, manifold signal reconstruction, and climate data imputation. APS significantly outperforms existing approaches in computational efficiency while maintaining competitive reconstruction quality, enabling real-time, high-quality solutions to inverse problems across diverse domains.
@@ 1 Introduction
${\bm{y}}={\mathcal{A}}({\bm{x}})+{\bm{n}},\quad{\bm{y}}\in{\mathbb{R}}^{m},\,{\bm{x}}\in{\mathbb{R}}^{n},\,{\mathcal{A}}:{\mathbb{R}}^{n}\mapsto{\mathbb{R}}^{m},\,{\bm{n}}\sim{\mathcal{N}}(0,\sigma_{y}^{2}{\bm{I}}),$ (1)
where the goal is to infer an unknown signal ${\bm{x}}$ from the degraded measurement ${\bm{y}}$ obtained through some forward operator ${\mathcal{A}}$, leveraging the information contained in the measurement and prior $p({\bm{x}})$. A powerful modern way to define the prior is through diffusion models (Ho et al., 2020; Song et al., 2021c), where we train parametrized model ${\bm{s}}_{\theta}$ to estimate the gradient of log prior $\nabla_{\bm{x}}\log p({\bm{x}})$.
Solving inverse problems with the diffusion model can be achieved through posterior sampling with Bayesian inference. Arguably the standard way to achieve this is through modifying the reverse diffusion process of diffusion models (Daras et al., 2024). This adjustment shifts the focus from sampling from the trained prior distribution $p_{\theta}({\bm{x}}_{0})$ to sampling from the posterior distribution $p_{\theta}({\bm{x}}_{0}|{\bm{y}})$. This transition is facilitated by employing iterative projections to the measurement subspace (Kadkhodaie & Simoncelli, 2021; Song et al., 2021c; Chung et al., 2022b; Wang et al., 2023), guiding the samples through gradients pointing towards measurement consistency (Chung et al., 2023a; Song et al., 2023a).
It should be noted that diffusion models learn the gradient of the prior and diffusion samplers (Song et al., 2021a; Lu et al., 2022; Song et al., 2021c) are methods that numerically solve the probability-flow ODE (PF-ODE) that defines the reverse diffusion sampling trajectory. Consequently, regardless of the specifics of the methods, standard diffusion model-based inverse problem solvers (DIS), even those that are considered fast, take at least a few tens of neural function evaluation (NFE), making them less effective for time-critical applications such as medical imaging and computational photography.
@@
Another class of methods (Feng et al., 2023; Feng & Bouman, 2024) introduces the use of variational inference (VI) to train a new proposal distribution $q_{\phi}^{\bm{y}}({\bm{x}})$ to distill the prior learned through the pre-trained diffusion model. The problem is defined as the following optimization problem
@@
where the superscript ${\bm{y}}$ emphasizes that the proposal distribution is specific for a single measurement ${\bm{y}}$. For tractable optimization, $q$ is often taken to be a normalizing flow (Rezende & Mohamed, 2015; Dinh et al., 2016) (NF) so that the computation of likelihood can be done instantly, and one can sample multiple reconstructions from the posterior samples by plugging in different noise values from the reference distribution. Posing the problem this way yields a method that can achieve posterior samples with just a single NFE. Nevertheless, it is still impractical as training a measurement-specific variational distribution takes hours of training. It is often unrealistic to train a whole new model from scratch every time when a new measurement is taken.
Figure 2: Concept of the proposed method, APS. (a) Training can be performed in an unsupervised fashion with a dataset consisting of degraded measurements ${\bm{y}}$ to train a conditional normalizing flow $G_{\phi}$ with the diffusion prior ${\bm{s}}_{\theta}$. (b) Once trained, one can achieve multiple posterior samples by inputting different noise vectors ${\bm{z}}\sim{\mathcal{N}}(0,{\bm{I}})$ concatenated with the condition ${\bm{y}}$ with a single NFE, generalizable across any measurement ${\bm{y}}$.
In this work, we take a step towards a practical VI-based posterior sampler by distilling a diffusion model prior. To this end, we propose a conditional normalizing flow $q_{\phi}({\bm{x}}_{0}|{\bm{y}})$ as our variational distribution and amortize the optimization problem in Eq. (2) over the conditions ${\bm{y}}$. By using a network that additionally takes in the condition ${\bm{y}}$ as the input, we can train a single model that generalizes across the whole dataset without the need for cumbersome re-training for specific measurements. (See Fig. 2 for the conceptual illustration of the proposed method, as well as representative results presented in Fig. 1.) Interestingly, we find that the speed of optimization is not hampered with such amortization, and the proposed method achieves comparable performance against the measurement-specific flow model (Feng & Bouman, 2024; Feng et al., 2023). Furthermore, we extend the theory to consider inverse problems on the Riemannian manifold, showing that the proposed idea is generalizable even when the signal is not on the Euclidean manifold. In summary, our contributions and key takeaways are as follows
@@
    To the best of our knowledge, our method is the first diffusion prior distillation approach for solving inverse problems that are unsupervised (i.e. does not require any ground-truth data ${\bm{x}}$), as opposed to standard conditional NFs (Lugmayr et al., 2020) that required supervised paired data.
@@ 2.1 Score-based diffusion models
We adopt the standard framework for constructing a continuous diffusion process ${\bm{x}}(t)$, where $t\in[0,T]$ and ${\bm{x}}(t)\in\mathbb{R}^{d}$, as outlined by song2020score. Specifically, our goal is to initialize ${\bm{x}}(0)$ from a distribution $p_{0}({\bm{x}})=p_{\text{data}}$, and evolve ${\bm{x}}(t)$ towards a reference distribution $p_{T}$ at time $T$, which is easy to sample from.
@@
where $dt$ represents an infinitesimal negative time step, and $\bar{{\bm{w}}}$ denotes the backward standard Brownian motion. While the trajectory of Eq. (4) is stochastic, there also exists a corresponding probability-flow ODE (PF-ODE) that recovers the same law $p_{t}({\bm{x}})$ as the time progresses (song2020score; song2020denoising)
@@
This allows a deterministic mapping between the reference and the target distribution, and hence diffusion models can also be seen as a neural ODE (chen2018neural).
A neural network can be trained to approximate the true score function $\nabla_{\bm{x}}\log p_{t}({\bm{x}})$ through score matching techniques, as demonstrated in previous works (song2019generative; song2020score). This approximation, denoted ${\bm{s}}_{\theta}({\bm{x}},t)\approx\nabla_{\bm{x}}\log p_{t}({\bm{x}})$, is then utilized to numerically integrate the reverse-time SDE. To effectively train the score function, denoising score matching (DSM) is often employed (hyvarinen2005estimation)
$\theta^{*}=\operatorname*{arg\,min}_{\theta}{\mathbb{E}}_{t\sim U(\varepsilon,1),{\bm{x}}(t),{\bm{x}}(0)}\left[\|{\bm{s}}_{\theta}({\bm{x}}(t),t)-\nabla_{{\bm{x}}_{t}}\log p_{0t}({\bm{x}}(t)|{\bm{x}}(0))\|_{2}^{2}\right],$ (6)
Interestingly, the posterior mean, or the so-called denoised estimate can be computed via Tweedie’s formula (efron2011tweedie). Specifically, for $p({\bm{x}}_{t}|{\bm{x}}_{0})={\mathcal{N}}({\bm{x}}_{t};\alpha_{t}{\bm{x}}_{0},\beta_{t}^{2}{\bm{I}})$,
@@ 2.2 Diffusion models for inverse problems (DIS)
While the former term can be replaced with a pre-trained diffusion model, the latter term is intractable and needs some form of approximation. Existing DIS (kawar2022denoising; chung2023diffusion; wang2023zeroshot) propose different approximations for $\nabla_{{\bm{x}}_{t}}\log p({\bm{y}}|{\bm{x}}_{t})$, which yields sampling from slightly different posteriors $\nabla_{{\bm{x}}_{t}}\log p_{\theta}({\bm{x}}_{t}|{\bm{y}})$.
Algorithmically, the posterior samplers are often implemented so that the original numerical solver for sampling from the prior distribution remains intact, while modifying the Tweedie denoised estimate at each time $\hat{\bm{x}}_{0|t}^{\theta}$ to satisfy the measurement condition given as Eq. (1). From Tweedie’s formula, we can see that this corresponds to approximating the conditional posterior mean ${\mathbb{E}}[{\bm{x}}_{0}|{\bm{x}}_{t},{\bm{y}}]$ in the place of the unconditional counterpart ${\mathbb{E}}[{\bm{x}}_{0}|{\bm{x}}_{t}]$. The algorithms are inherently iterative, and the modern solvers (chung2023diffusion; wang2023zeroshot; zhu2023denoising) require at least 50 NFE to yield a high-quality sample. Moreover, as existing methods can be interpreted as approximating the reverse distribution $p({\bm{x}}_{0}|{\bm{x}}_{t})$ with a simplistic Gaussian distribution $q({\bm{x}}_{0}|{\bm{x}}_{t})={\mathcal{N}}({\bm{x}}_{0};\hat{\bm{x}}_{0|t}^{\theta},s_{t}^{2}{\bm{I}})$ (peng2024improving), it typically yields a large approximation error, especially in the earlier steps of the reverse diffusion.
@@
Methods DIS Noise2Score RED-Diff Score prior APS (ours)
@@ 3.1 Variational inference in DIS
Standard DIS discussed in Sec. 2.2 sample from the posterior distribution by following the reverse diffusion trajectory. Another less studied approach uses VI to use a new proposal distribution, where the problem is cast as an optimization problem in Eq. (2). RED-diff (mardani2023variational) places a unimodal Gaussian distribution as the proposal distribution $q_{\phi}^{\bm{y}}({\bm{x}})$, and the KL minimization is done in a coarse-to-fine manner, similar to standard DIS, starting from high noise level to low noise level. While motivated differently, RED-diff and standard DIS have similar downsides of requiring at least a few tens of NFEs, as well as placing a simplistic proposal distribution. Furthermore, one can achieve only a single sample per optimization.
Recently, Feng et al. (feng2023score; feng2024variational) uses an NF model for the proposal distribution while solving the same VI problem. The optimization problem involves computing the diffusion prior log likelihood $\log p_{\theta}({\bm{x}})$. It was shown that it can be exactly computed by solving the PF-ODE (feng2023score; song2020score), but numerically solving the PF-ODE per every optimization step is extremely computationally heavy, and hence does not scale well. To circumvent this issue, it was proposed to use a lower bound (feng2024variational; song2021maximum). Once trained, the NF model can be given different noise inputs ${\bm{z}}\sim{\mathcal{N}}(0,{\bm{I}})$ to generate diverse posterior samples with a single forward pass through the network. However, the training should be performed with respect to all the different measurements, not being able to generalize across the dataset. Our work follows along this path to overcome the current drawback and optimize a single model for entire measurement space with a similar cost as shown in Tab. 1.
@@ 3.2 Distillation of the diffusion prior
Our method involves distillation of the diffusion prior into a student deep neural network, in our case an NF model. Particularly, it involves evaluating the output of the model by checking the denoising loss gradients from the pre-trained diffusion model. This idea is closely related to variants of score distillation sampling (SDS) (poole2023dreamfusion; wang2024prolificdreamer), where the gradient from the denoising loss is used to distill the diffusion prior by discarding the score Jacobian. Possibly a closely related work is Diff-instruct (luo2024diff), where the authors propose to train a one-step generative model similar to GANs (goodfellow2014generative) by distillation of the diffusion prior with VI. By proposing an integral KL divergence (IKL) by considering KL minimization across multiple noise levels across the diffusion, it was shown that SDS-like gradients can be used to effectively train a new generative model. While having similarities, our method directly minimizes the KL divergence and does not require dropping the score Jacobian.
Orthogonal to the score distillation approaches, there have been recent efforts to train a student network to emulate the PF-ODE trajectory itself (song2023consistency; gu2023boot) with a single NFE, one of the most prominent directions being consistency distillation (CD) (song2023consistency). While promising, the performance of CM is upper-bounded by the teacher PF-ODE. Thus, in order to leverage CD-type approaches for diffusion posterior sampling, one has to choose one of the approximations of DIS as its teacher model. In this regard, applying CD for diffusion inverse problem solving is inherently limited.
4 Amortized Posterior Sampling (APS)
@@ 4.1 Conditional NF for amortized score prior
The goal is to use a variational distribution that is conditioned on ${\bm{y}}$, such that the resulting distilled conditional NF model $G_{\phi}$ generalizes to any condition ${\bm{y}}$. To this end, inspired from the choices of (sun2021deep; feng2023score) we modify the objective in Eq. (2) to
@@
Computation of $\log p_{\theta}(\hat{\bm{x}}_{0})$ is more involved: to exactly compute the value, we would have to solve the PF-ODE, which is compute-heavy (song2020score; feng2023score). To circumvent this burden, we leverage the evidence lower bound (ELBO) (song2021maximum; feng2024variational) $b_{\theta}(\hat{\bm{x}}_{0})\leq\log p_{\theta}(\hat{\bm{x}}_{0})$: